OpenStudy (anonymous):

MN = 9 Root 2 LM= ? (pic) PLEASE HELP

OpenStudy (anonymous):



OpenStudy (anonymous):

its a 30 60 90 triangle

OpenStudy (anonymous):

ln is twice of the leg

OpenStudy (anonymous):

what is that \[7\sqrt{2}\]???

OpenStudy (anonymous):

@SarahNicole

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[7 \sqrt{2}\]

OpenStudy (anonymous):

im not to sure how to do that

OpenStudy (anonymous):

14\[\sqrt{2}\]

OpenStudy (anonymous):

\[14\sqrt{2}\]

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

yup you are correct like I said hypotenuse is twice of leg

OpenStudy (anonymous):

i have one more like that would u like to help

OpenStudy (anonymous):

@SarahNicole just multiply the leg by two to get the hypotenuse it applies 30-60-90 triangle

OpenStudy (anonymous):

the same triangle just mn=\[9\sqrt{2}\] find LM

OpenStudy (anonymous):